OpenStudy (anonymous):

WILL GIVE MEDALS A box company completed a study to determine the number of boxes an old machine would produce in an hour. They determined that the old machine produced 280 boxes in 4 hours. During its first day of operation, a new machine, after running a full 8 hours, produced 56 more boxes than the old machine would have produced in that time. Which of the following is true about the new machine?

OpenStudy (anonymous):

The new machine can produce 14 boxes an hour. The equation y = 84x represents the production rate of the new machine. The new machine can produce 7 more boxes an hour than the old machine. The equation y = 77x represents the production rate of the new machine.

OpenStudy (anonymous):

need two choose two

OpenStudy (anonymous):

which question

OpenStudy (anonymous):

A box company completed a study to determine the number of boxes an old machine would produce in an hour. They determined that the old machine produced 280 boxes in 4 hours. During its first day of operation, a new machine, after running a full 8 hours, produced 56 more boxes than the old machine would have produced in that time. Which of the following is true about the new machine? During its first day of operation, a new machine, after running a full 8 hours, produced 56 more boxes than the old machine would have produced in that time. Which of the following is true about the new machine? A box company completed a study to determine the number of boxes an old machine would produce in an hour. They determined that the old machine produced 280 boxes in 4 hours. During its first day of operation, a new machine, after running a full 8 hours, produced 56 more boxes than the old machine would have produced in that time. Which of the following is true about the new machine?

OpenStudy (anonymous):

oh okay i get it thanks

OpenStudy (anonymous):

SORRY its all the same pAST to much

OpenStudy (anonymous):

oh the PSAT?

OpenStudy (anonymous):

can u put the answer choices like separate so i can see just to make it easier thanks.

OpenStudy (anonymous):

um....

OpenStudy (anonymous):

old machine 280/4 = 70 per hour or 560 for 8 hours new machine = 560 + 56 = 616 boxes in 8 hours

OpenStudy (anonymous):

new machine 616/8 = 77 boxes per hour (hint - one of your choices)

OpenStudy (anonymous):

Does that make sense?